OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 57 miles at a certain speed. The canoeist travels 5 miles on the second part of the trip at a speed 5 mph slower. The total time of the trip is 3 hours. what was the speed on each part of the trip

OpenStudy (anonymous):

i think you should try it first

OpenStudy (wasiqss):

first journey speed is 57/x second journey speed (5/x)-5 equate time to get answer

OpenStudy (dumbcow):

distance = rate*time Let x be original speed, t be time for 1st part of trip 57 = xt 5 = (x-5)(3-t) solve system of 2 equations

OpenStudy (dumbcow):

use substitution: t = 57/x 5 = (x-5)(3-57/x) Now multiply both sides by x to get it out of denominator 5x = (x-5)(3x -57) Distribute 5x = 3x^2-72x +285 set equal to 0 3x^2 -77x +285 = 0 use quadratic formula x = 77 +-sqrt(77^2 - 4(3)(285)) / 6 x = 77 +- sqrt(2509)/6 x = 4.485 or 21.18 x must be greater than 5, since x-5 is speed on 2nd part of trip x= 21.18 x-5 = 16.18